Sreening effect

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The screening effect, or shielding effect, is how electrons in the same atom interact with each other. In a single-electron atom (in isolation), the electron is only interacting with the proton; in a multielectron atom, the electrons are both interacting with the proton(s), but also with each other. While electrons are attracted to the protons in the nucleus, they are repelled by the other electrons. This electron-electron repulsion decreases the attractive force of the protons on the electrons.

The shielding effect changes the effective nuclear charge -- effectively decreasing the true nuclear charge. This effect causes atoms to get smaller as you across a period (row) of the periodic table, as well as many other periodic trends observed in the periodic table.

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vijesh jayan
13 years ago

In solids, especially in metals and semiconductors, the electrostatic screening of screening effect reduces the electrostatic field and Coulomb potential of an ion inside the solid. Like the electric field of the nucleus is reduced inside an atom or ion due to the shielding effect, the electric fields of ions in conducting solids are further reduced by the cloud of conduction electrons. The screened Coulomb potential is expressed as

V(r) = \frac{Z e}{r} e^{-q r}

where Z is the atomic number, e is the elementary unit charge, r is the distance to the nucleus of the embedded ion and q is a screening parameter that determines the range of the potential. The screening parameter q plays an important role in theoretical models in solid-state physics  The screened electrostatic potential, like the Yukawa potential, has a simple Fourier transform, expressed as

V(k) = \frac{4 \pi Z e}{q^2 + k^2}

The screened potential determines the inter atomic force and the phonon dispersion relation in metals. The screened potential is used to calculate the electronic band structure of a large variety of materials, often in combination with pseudopotential models,

 

APPROVE PLZZZZZZZZZZZZZZZ